Bilbo Baggins wants to save money to meet three objectives. First, he would like to be able to retire 30 years from now with retirement income of $28,000 per month for 30 years, with the first payment received 30 years and 1 month from now. Second, he would like to purchase a cabin in Rivendell in 20 years at an estimated cost of $1,239,000. Third, after he passes on at the end of the 30 years of withdrawals, he would like to leave an inheritance of $550,000 to his nephew Frodo. He can afford to save $1,700 per month for the next 20 years. If he can earn a 12 percent EAR before he retires and a 11 percent EAR after he retires, how much will he have to save each month in Years 21 through 30?

$12,524.57
$9,774.14
$9,390.84
$9,582.49
$12,450.30
Answer #1

PE EPS, STOCK DIVIDEND - Microsoft Excel (Product Activation Failed) View Add-Ins Review - 2x = = Dr General Wrap Text Merge & Center Σ AutoSum : A Fill 2 Sort & to be conditional Format set $ - % m Find & 8. inser Formatting as Table Styles Styles Clear Select Filter Editing Alignment Number Cells F G H M File Home Insert Page Layout Formulas Data % Cut Calibri 11AA = Ea Copy Paste Format Painter B I U D A E Clipboard Font G346 A B 331 332 STEP 1 PRE RETIREMENT APR 333 EAR 334 NOMINAL RATE 335 336 STEP 2 POST RETIREMENT APR 337 EAR 338 NOMINAL RATE 12% 11.3866% EXCEL =NOMINAL(12%,12) 11% 10.4815% EXCEL = NOMINAL(11%, 12) 339 STEP 3 340 341 PV OF 28000 PER MONTH PV OF 550000 (inheritance) 3065612.63 24025.55 [=PV(D338/12,12*30,-28000)] [=PV(D338/12,12*30,,-550000)] 342 SO HE NEEDS AT RETIREMENT 3089638.18 [=SUM(D340:D341)] 343 344 345 STEP 4 346 347 348 349 SAVINGS RATE YEARS EV OF SAVINGS 1700 11.3866% 20 1549058.81 [=FV(D334/12,D347*12,-D345)] N d marginal tax PB ROPI APV ROSS price quote EVA MVA LEVERAGED BUY SWAP distress BRIGHAN CAP GAIN Sheet1 I eberk MARGIN MONEY ferences call put bilbo NI BO. 130% Rea 40-_+ EJER * ENG 25-09-2019